# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

### v1

@@ VIII Eikonal approximation for HP contribution
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})(q_{t}^{2})^{-\epsilon/2}\frac{64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}}{q_{\perp}^{2}}\Big{\{}\frac{2}{2-\epsilon}+(1-\frac{2}{2-\epsilon})\Big{\}}.$ (202)
Now the two $2/(2-\epsilon)$ cancel out. From above derivation we see clearly how the cancellation occurs. The divergence of $I\langle L\rangle$ is given by
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}\frac{(q_{t}^{2})^{-\epsilon/2}}{q_{\perp}^{2}}$
$=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\frac{-64\cos\theta l_{\perp,cs}\cdot\tilde{s}_{\perp}}{Q^{2}}\Big{(}\frac{4}{\epsilon^{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})+\cdots\Big{)}.$ (203)
where $\cdots$ in brackets does not contain the two delta functions at the same time.
Above formulas for $I\langle L\rangle$ also hold for $I\langle P_{7}\rangle$. But because $P_{7}\propto q_{\perp}$, only the derivative term in $N_{2}^{\rho}|_{\text{II}}$ contributes. Because
$g_{\perp\mu\nu}\frac{\partial P_{7}^{\mu\nu}}{\partial q_{\perp}^{\rho}}=\frac{1}{\tilde{s}_{\perp}^{2}}g_{\perp\mu\nu}(g_{\perp}^{\mu\rho}\tilde{s}_{\perp}^{\nu}+g_{\perp}^{\nu\rho}\tilde{s}_{\perp}^{\mu})=\frac{2\tilde{s}_{\perp}^{\rho}}{\tilde{s}_{\perp}^{2}},$ (204)
we have
$I\langle P_{7}\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\frac{(q_{t}^{2})^{-\epsilon/2}}{q_{\perp}^{2}}\frac{-32}{2-\epsilon}.$ (205)
The factor $1/(2-\epsilon)$ is left. Now, it is clear that under eikonal approximation, the divergent part of $I\langle L\rangle$ and $I\langle P_{7}\rangle$ satisfies following relation
@@
$I\langle L\rangle=-4(1-\frac{\epsilon}{2})\tilde{s}_{\perp}\cdot l_{\perp,cs}\cos\theta I\langle P_{7}\rangle,$ (206)
which is consistent with our previous explicit results eqs.(127,128).
Moreover, if there is no the second line in eq.(200), we have $I\langle L\rangle=-4I\langle P_{7}\rangle$, then the factorization holds. The tensor in the second line of eq.(200),
$(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})$ (207)
has a coefficient proportional to
$(1-\frac{2}{2-\epsilon})=\frac{-\epsilon}{2-\epsilon}.$ (208)
But because $1/q_{\perp}^{2}$ gives $1/\epsilon^{2}$, this tensor still has a divergent coefficient. It is this coefficient that cannot be subtracted. The same tensor also appears in virtual correction. See eq.(91), where
$g_{3,NP}^{\mu\nu}=C_{3,NP}t_{a}^{\mu\nu}+D_{3,NP}t_{b}^{\mu\nu}=C_{3,NP}\frac{1}{x_{2}}\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}.$ (209)
However, $g_{3,NP}^{\mu\nu}$ has only non-pole contribution($x_{1}\neq x_{2}$), which cannot cancel the divergence given here. Thus, the soft divergence associated with above tensor in real correction cannot be cancelled by corresponding virtual correction. Moreover, from eq.(191), the gluon with momentum $k_{g}-k$ behaviors like $_{g}^{+}-k^{+}\sim\lambda^{2}, kk$$_{\perp}-k_{g\perp}\sim\lambda$($k_{\perp}=0$). So, $|(k_{g}^{+}-k^{+})(k_{g}^{-}-k^{-})|\ll|k_{g\perp}-k_{\perp}|^{2}$. The gluon thus is a Glauber gluon. In unpolarized DY, the Glauber gluon contribution is shown to cancel out, see for example Bodwin (1985). Our above analysis indicates that such a cancellation does not happen for SSA.
IX Summary
@@
In this work, we have calculated the one-loop correction to unweighted SSA in DY for the lepton angular distribution. As a comparison, we also calculate a weighted observable $I\langle P_{7}\rangle$ with the same method. We use at most one $G^{+}$ to do collinear expansion, and then use EOM for fermion to eliminate the bad component of fermion field. After this, the hadronic tensor or cross section is expressed by independent twist-3 distribution functions. Three of these functions are $q_{\partial}$, $T_{F}$ and $T_{\Delta}$. For virtual correction, we first determine that the hard part is analytic on the upper half planes of $s_{0},s_{1}$. After this is clear, we use FIRE to reduce the various tensor integrals in the hard part to standard scalar integrals. Due to the simple kinematics, the reduced scalar integrals are three bubble integrals $B(s_{0})$, $B(s_{1})$ and $B(s_{2})$. Then, the structure of virtual correction becomes clear. We successfully extract pole part and non-pole part. For pole part, only SGP is possible. QCD and QED gauge invariance are maintained in the calculation. On the other hand, for real correction, we first assume $q_{\perp}\neq 0$. Then, the tensor structure of $W^{\mu\nu}$ is worked out. With the projection tensors we calculate the real corrections to $I\langle L\rangle$ in the same way as $I\langle P_{7}\rangle$. Only three-point distribution functions contribute at one-loop level. The contributions are classified into SGP, HP and SFP contributions. Corresponding contributions are worked out explicitly. In this work, only contributions from $\bar{q}(x_{b})$ and $T_{F}(x_{1},x_{2})$ are retained. For both virtual and real corrections, $I\langle P_{7}\rangle$ is the same as our previous resultChen et al. (2017). Then, the collinear subtraction is performed. We find all divergences in $I\langle P_{7}\rangle$ can be subtracted out, while a single pole in $I\langle L\rangle$ cannot be subtracted. The single pole is proportional to $N_{c}C_{A}/\epsilon$ and $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$. In addition, we find the divergent nonpole part of virtual correction can be viewed as a part of HP contribution of real correction, and can be subtracted without any problem. The only term to break the factorization is the single pole term we mentioned above. We also analyze the source for this unsubtracted divergence under eikonal approximation for real HP corrections. We find in $W^{\mu\nu}$ a tensor appears with a soft divergence, but such a tensor in virtual correction has no corresponding soft divergence. We thus conclude that it is this uncancelled soft divergence that breaks the factorization. Very recently, the one-loop correction to SSA in lepton-hadron scattering is calculated in Rein et al. (2025a, b). No breaking of factorization is found. In future, we will study this process using present method. Moreover, the breaking of factorization found here depends on collinear expansion. We would like to check this conclusion in future with the multiparton state proposed in Cao et al. (2010); Ma et al. (2012).
@@ Acknowledgements
The author would like to thank J.P. Ma for very helpful discussions. The hospitality of ITP, CAS(Beijing) is appreciated during the finalization of this paper.
@@ Appendix A Counter terms of QCD
In $\overline{MS}$ scheme, the values of these constants are well known, and can be found in e.g.,Muta (1998). For convenience, we list the values in Feynman gauge as follows:
@@
Since Fig.5(c) causes the problem, we give the details for this diagram. For hard pole, $k^{+}\neq 0$, we can use $G_{\perp}$ to calculate directly. With $G_{\perp}$ as the initial gluon, the hard part is
@@
$I\langle L\rangle=$ $\tilde{s}_{\perp}^{\rho}\int d^{n}q\delta(q^{2}-Q^{2})\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2}).$ (218)
@@
$I\langle L\rangle=$ $\tilde{s}_{\perp}^{\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\int d^{n-2}q_{\perp}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2})$
$=$ $\tilde{s}_{\perp}^{\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\frac{\Omega_{n-2}}{2}(q_{t}^{2})^{-\epsilon/2}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\hat{x}_{2}.$ (220)

### v2

@@ VII.3 Final results
$I\langle L\rangle|_{final}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int\frac{d\xi}{\xi}\Big\{-N_{c}\ln\frac{\mu^{2}}{Q^{2}}\Big[\Big(P_{qq}\otimes\bar{q}(x_{b}^{*})\Big)T_{F}(x_{2}^{*},x_{2}^{*})+\bar{q}(x_{b}^{*})\Big(\mathcal{F}_{q}\otimes T_{F}(x_{2}^{*})\Big)\Big]$
$+\bar{q}(x_{b}^{*})T_{F}(x_{2}^{*},x_{2}^{*})(N_{c}^{2}-1)(\frac{\pi^{2}}{2}-4)+\bar{q}(x_{b}^{*})\int_{x_{2}^{*}-1}^{1}dx_{2}T_{F}(x_{2}^{*}-x_{2},x_{2}^{*})\Big[N_{c}^{2}(2+\ln x_{2}^{*})\delta(x_{2})$
$+\frac{2x_{2}^{*}}{x_{2}^{2}}\theta(x_{2}-x_{2}^{*})+\frac{N_{c}^{2}\theta(-x_{2})}{x_{2}^{*}-x_{2}}-\frac{N_{c}^{2}\theta(x_{2})}{(x_{2})_{+}}+\frac{1}{(x_{2})_{p}}+\frac{x_{2}^{*}\theta(x_{2}^{*}-x_{2})}{(x_{2}^{2})_{p}}\ln\frac{x_{2}^{*}-x_{2}}{x_{2}^{*}}\Big]\Big\}$
$+I\langle L\rangle\Big|_{r}^{SGP}+I\langle L\rangle\Big|_{r}^{HP-1}+I\langle L\rangle\Big|_{r}^{HP-2}+I\langle L\rangle\Big|_{r}^{HP-3}+I\langle L\rangle\Big|_{r}^{SFP}.$ (191)
The last line is the reminder part of real corrections after subtraction. Their explicit expressions are given below.
SGP part is:
@@
$I\langle L\rangle|_{r}^{SGP}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int_{r}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big[g_{1}^{(a)}\delta(1-\hat{x}_{b})+g_{2}^{(a)}\delta(1-\hat{x}_{2})+\frac{g_{3}^{(a)}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big],$
$g_{1}^{(a)}=$ $-\frac{1}{2}\Big[L_{2}(1+\hat{x}_{2}^{2})-\hat{x}_{2}+\hat{x}_{2}^{2}\Big],$
$g_{2}^{(a)}=$ $-\frac{1}{2}\Big[L_{1}\hat{x}_{b}(1+\hat{x}_{b}^{2})+\hat{x}_{b}-1\Big],$
$g_{3}^{(a)}=$ $\frac{\hat{x}_{2}+\hat{z}-1}{2Q^{2}\hat{x}_{2}\hat{z}^{2}}\Big[Q\hat{x}_{2}^{3}\left(\hat{z}-1\right)+\hat{x}_{2}^{2}\left(\hat{z}\left(E_{t}-Q\right)+Q\right)+\hat{x}_{2}\left(\hat{z}-1\right)\hat{z}\left(E_{t}+Q\hat{z}\right)+Q\hat{z}^{2}\left(\hat{z}^{2}-\hat{z}+1\right)\Big].$ (192)
There are three parts of hard pole contributions, which correspond to different $T_{F}$. The first part is:
$I\langle L\rangle|_{r}^{HP-1}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\Big[g_{1}^{(b)}\delta(1-\hat{x}_{b})+g_{2}^{(b)}\delta(1-\hat{x}_{2})+\frac{g_{3}^{(b)}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big],$
$g_{1}^{(b)}=$ $\frac{N_{c}^{2}}{2}\Big[L_{2}(1+\hat{x}_{2})-1\Big],$
$g_{2}^{(b)}=$ $\frac{\hat{x}_{b}+N_{c}^{2}-1}{2}\Big[L_{1}(1+\hat{x}_{b}^{2})+2\Big],$
$g_{3}^{(b)}=$ $\frac{N_{c}^{2}+\hat{z}-1}{2\hat{x}_{2}E_{t}}\Big[\hat{x}_{2}\hat{z}E_{t}+\left(\hat{z}-1\right)\hat{z}E_{t}+Q\hat{x}_{2}^{2}\Big];$ (193)
The second part is:
$I\langle L\rangle|_{r}^{HP-2}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\Big[g_{1}^{(c)}\delta(1-\hat{x}_{b})+\frac{g_{3}^{(c)}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}\Big],$
$g_{1}^{(c)}=$ $\frac{1}{2}\Big[\ln(1-\hat{x}_{2})-1\Big],$
$g_{3}^{(c)}=$ $-\frac{Q^{3}\hat{x}_{2}(\hat{x}_{2}-1)(\hat{z}-1)(2N_{c}(\hat{z}-1)+\hat{z})}{2E_{t}^{3}(\hat{x}_{2}+\hat{z}-1)}+\frac{(\hat{x}_{2}-1)(\hat{x}_{2}+\hat{z}-1)}{2\hat{x}_{2}\hat{z}}\Big[N_{c}(\hat{z}-1)(\hat{z}-2)-\hat{z}\Big$]. (194)
The third part is:
$I\langle L\rangle|_{r}^{HP-3}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int_{r}\bar{q}(x_{b})T_{F}(x_{2}-x_{1}^{*},-x_{1}^{*})\frac{g_{3}^{(d)}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}},$
$g_{3}^{(d)}=$ $-\frac{(\hat{x}_{2}-1)(\hat{z}-1)(\hat{x}_{2}+\hat{z}-1)}{2\hat{x}_{2}\hat{z}}\Big[N_{c}(\hat{z}-2)-\hat{z}+\hat{z}^{2}\Big]+\frac{Q^{3}\hat{x}_{2}(\hat{x}_{2}-1)(\hat{z}-1)^{2}(2N_{c}+\hat{z})}{2E_{t}^{3}(\hat{x}_{2}+\hat{z}-1)}.$ (195)
The last part, SFP contribution, is:
$I\langle L\rangle|_{r}^{SFP}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{\pi N_{c}^{2}}\int_{r}\bar{q}(x_{b})T_{F}(0,-x_{2})\frac{g_{3}^{(e)}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}},$
$g_{3}^{(e)}=$ $-\frac{(\hat{x}_{2}-1)(\hat{z}-1)(\hat{x}_{2}+\hat{z}-1)}{2Q\hat{x}_{2}\hat{z}}\Big[\hat{x}_{2}E_{t}+2Q\hat{x}_{2}^{2}-Q\hat{z}^{2}\Big].$ (196)
Note that all explicit $\ln\mu$ dependence is provided by subtraction in our formalism. With tree level result taken into account, our result is $\mu$ independent to $O(\alpha_{s})$.
VIII Summary
@@
In this work, we have calculated the one-loop correction to unweighted SSA in DY for the lepton angular distribution. We choose to work in Feynman gauge. As a comparison, we also calculate a weighted observable $I\langle P_{7}\rangle$ with the same method. We use at most one $G^{+}$ to do collinear expansion, and then use EOM for fermion to eliminate the bad component of fermion field. After this, the hadronic tensor or cross section is expressed by independent twist-3 distribution functions. Three of these functions are $q_{\partial}$, $T_{F}$ and $T_{\Delta}$, which are gauge invariant. The other two distribution functions are not gauge invariant and their hard coefficients are found to be zero. To calculate virtual correction, we first illustrate that the hard part is analytic on the upper half planes of $s_{0},s_{1}$. After this we use FIRE to reduce the various tensor integrals to standard scalar integrals. Due to the simple kinematics, the reduced scalar integrals are three bubble integrals $B(s_{0})$, $B(s_{1})$ and $B(s_{2})$. According to the analyticity of the amplitude on $s_{0},s_{1}$, we successfully extract pole part and non-pole part of virtual corrections. Both are divergent, and for the pole part only soft gluon pole contributes. Virtual corrections to $W^{\mu\nu}$ can be written into a form consistent with QED gauge invariance. For real corrections, only three kinds of pole contributions are possible, that is, SGP, HP and SFP contributions. There is no contribution from two-point twist-3 distribution functions. Then, we perform collinear subtraction. As found in [20, 21], the divergence of nonpole part of virtual correction actually can be written into the same form as hard pole contribution of real corrections. In the sum of virtual and real corrections, all divergences can be subtracted out by the one-loop corrections to twist-2 and twist-3 distribution functions. Thus, the twist-3 factorization still holds for polarized DY at one-loop level. We further give all subtracted finite hard coefficients. Only the coefficients of the convolution $\bar{q}\otimes T_{F}$ are considered in this work. In future, we will also extend the calculation to include chiral-odd and pure gluon twist-3 distribution functions. Moreover, some NLO corrections to SSA in lepton-hadron scattering are worked out in [20, 21]. The factorization formula there has a more complicated form, e.g., the derivative of $T_{F}(x,x)$ appears. We will consider this process in future to check our formalism further.
@@ Acknowledgements
The author would like to thank Prof. Jian-Ping Ma for very helpful discussions. The hospitality of ITP, CAS(Beijing) during the finalization of this paper is appreciated.
@@ Appendix A Counter terms of QCD
In $\overline{MS}$ scheme, the values of these constants are well known, and can be found in e.g.,[44]. For convenience, we list the values in Feynman gauge as follows:
@@
For hard pole, $k^{+}\neq 0$, we can use $G_{\perp}$ to calculate directly. With $G_{\perp}$ as the initial gluon, one of the hard parts of Fig.5(c) is
@@
$I\langle L\rangle=$ $\tilde{s}_{\perp\rho}\int d^{n}q\delta(q^{2}-Q^{2})\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2}).$ (205)
@@
$I\langle L\rangle=$ $\tilde{s}_{\perp\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\int d^{n-2}q_{\perp}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\delta((k_{2}+k_{b}-q)^{2})$
$=$ $\tilde{s}_{\perp\rho}\int\frac{d\xi}{2\xi}\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}l_{\perp,cs}^{\rho}\frac{\Omega_{n-2}}{2}(q_{t}^{2})^{-\epsilon/2}\bar{H}(q_{\perp}^{2},p_{a}\cdot q,p_{b}\cdot q)\hat{x}_{2}.$ (207)
@@
Appendix C $\gamma_{5}$ convention
We use HVBM scheme, where $\gamma_{5}$ is defined by $\gamma_{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$. In the trace it gives $Tr(\gamma_{5}\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}\gamma^{\tau})=-4i\epsilon^{\mu\nu\rho\tau}$, with $\epsilon^{0123}=+1$. At the same time, spin vectors $s_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$ are defined as 4-dim vectors. Following identity holds if all matrices are defined in 4-dim spacetime,
$\gamma^{\mu}\gamma^{\rho}\gamma^{\nu}=$ $S^{\mu\rho\nu\sigma}\gamma_{\sigma}-i\epsilon^{\mu\rho\nu\sigma}\gamma_{5}\gamma_{\sigma},$
$S^{\mu\rho\nu\sigma}=$ $g^{\mu\rho}g^{\nu\sigma}+g^{\mu\sigma}g^{\nu\rho}-g^{\mu\nu}g^{\rho\sigma}.$ (213)
With this identity, it is obvious that
[…]
Note that $\tilde{s}_{\perp}^{\mu}=\epsilon_{\perp}^{\mu\nu}s_{\perp\nu}$, $s_{\perp}^{\mu}=-\epsilon_{\perp}^{\mu\nu}\tilde{s}_{\perp\nu}$. This scheme is crucial for getting correct virtual corrections.